# TransKD: Transformer Knowledge Distillation for Efficient Semantic Segmentation

### v3

@@ Abstract
Semantic segmentation benchmarks in the realm of autonomous driving are dominated by large pre-trained transformers, yet their widespread adoption is impeded by substantial computational costs and prolonged training durations. To lift this constraint, we look at efficient semantic segmentation from a perspective of comprehensive knowledge distillation and consider to bridge the gap between multi-source knowledge extractions and transformer-specific patch embeddings. We put forward the Transformer-based Knowledge Distillation (TransKD) framework which learns compact student transformers by distilling both feature maps and patch embeddings of large teacher transformers, bypassing the long pre-training process and reducing the FLOPs by ${>}85.0\%$. Specifically, we propose two fundamental and two optimization modules: (1) Cross Selective Fusion (CSF) enables knowledge transfer between cross-stage features via channel attention and feature map distillation within hierarchical transformers; (2) Patch Embedding Alignment (PEA) performs dimensional transformation within the patchifying process to facilitate the patch embedding distillation; (3) Global-Local Context Mixer (GL-Mixer) extracts both global and local information of a representative embedding; (4) Embedding Assistant (EA) acts as an embedding method to seamlessly bridge teacher and student models with the teacher’s number of channels. Experiments on Cityscapes, ACDC, NYUv2, and Pascal VOC2012 datasets show that TransKD outperforms state-of-the-art distillation frameworks and rivals the time-consuming pre-training method. The source code is publicly available at https://github.com/RuipingL/TransKD.
@@ I Introduction
Semantic segmentation assigns category labels at a pixel-level (see examples in Fig. 1) and is a crucial tool in a wide range of applications, such as autonomous driving [1, 2, 3, 4] and navigation assistance for vulnerable road users [5, 6]. Driven by the achievements of deep learning, semantic segmentation has empowered precise scene comprehension in the context of autonomous driving, as evidenced by remarkable accuracies [7, 8]. However, there has been a notable oversight in addressing the computational limitations crucial for real-world applications, including the speed of inference and training, as well as memory footprint.
@@
While Convolutional Neural Networks (CNNs) have been the visual recognition frontrunners for almost a decade, they face problems capturing large-scale interactions due to their focus on local neighborhood operations and limited receptive fields. Vision transformers [7, 8, 13] overcome this challenge through the concept of self-attention [14], which learns to intensify or weaken portions of input at a global scale. This excellent quality of modeling long-range dependencies has led to transformers being on top of virtually all segmentation benchmarks [7, 8, 13]. Unfortunately, transformers are usually very large and don’t generalize well when trained on insufficient amounts of data due to the lack of inductive bias compared to CNNs [13], such as translation equivariance and locality. Thus, vision transformers often show less satisfactory performance on smaller task-specific datasets, and time-consuming pre-training process on a massive dataset is entailed [7, 13], e.g., ${\geq}100$ epochs on ImageNet [11]. The high computational and training costs constitute a significant bottleneck for real applications, where cumbersome models cannot be deployed on mobile devices, whereas smaller models suffer from inaccurate predictions [12, 15].
@@
A popular way to reduce the computational cost is knowledge distillation [17] which offers a mechanism for an effective knowledge transfer between a large teacher model and a smaller student network and has been widely studied for deep CNNs [12, 9, 10]. Gou et al. [16] distinguish three types of knowledge to be distilled (see Fig. 3a-c): response-based knowledge [12, 9, 17], feature-based knowledge [12, 18, 19], and relation-based knowledge [10, 20, 21]. The response-based and feature-based knowledge distillation utilize the output of a single specific layer of the teacher model (the last or the intermediate layer respectively), whereas the relation-based knowledge distillation tracks the relationship between multiple layers simultaneously. These paradigms are mainly designed with the CNN-to-CNN transfer in mind [12, 9, 10] and do not sufficiently cover knowledge distillation from transformer-specific intermediate blocks. The patch embedding method is a transformer-specific module to partition the input images or feature maps into a sequence of patches. According to the vanilla vision transformer [13], simply partitioning the input images or feature maps into large non-overlap patches is impractical to learn the spatial relations, which is complemented with optimized embedding methods, e.g. learnable positional embeddings [13], conditional positional embeddings [22, 23], shifted windows [24], and overlapping patch embeddings [25, 26]. Viewing that there is a large gap between the performances of the transformers with a simple patchifying process and optimized embedding methods, we recognize that the sequential information within patch embedding is significant and regard patch embeddings as a transformer-specific knowledge source.
@@
To achieve efficient semantic segmentation, we propose Transformer-based Knowledge Distillation (TransKD), the first framework for distilling knowledge from large-scale semantic segmentation transformers. We focus on a comprehensive distillation from different knowledge sources within the transformer-specific building blocks (our key idea is presented in Fig. 3d). TransKD distills the information from both (1) feature maps and (2) the transformer-specific patch embeddings. As mentioned before, patch embeddings differ from feature maps not only through the distinct shape and dimensionality but also in the inherent knowledge acquired through the patchifying process. We believe that extracting information not only from the feature maps but also from the transformer-specific patch embeddings harvests both, spatial and sequential relationships and is therefore important to maximize the potential of transformer-to-transformer knowledge distillation.
@@
Specifically, we design two fundamental and two optimization modules within the TransKD framework. (1) For feature map distillation, we build a relation-based scheme atop Knowledge Review [10] and introduce the Cross Selective Fusion (CSF) module to merge cross-stage feature maps via channel attention and construct the feature map distillation streams within hierarchical transformers. (2) For patch embedding distillation, we introduce a fundamental Patch Embedding Alignment (PEA) module to perform a dimensional transformation of patch embeddings along the channel dimension to facilitate multi-stage patch embedding distillation streams. (3) As known to the field, both neighboring and long-range relationships are vital for semantic segmentation [26, 27], our Global-Local Context Mixer (GL-Mixer) extracts both global and local information of a representative embedding for distillation. (4) Furthermore, considering that the large gap between the student’s and the teacher’s size negatively impacts the knowledge distillation results [28, 29], we present an Embedding Assistant (EA) module, which acts an embedder with teacher’s number of channels. EA helps to build a pseudo teacher assistant model by combining the student’s transformer blocks and seamlessly bridge teacher and student models.
@@
We demonstrate the benefits of our approach for real-world mobile applications on four benchmarks: a general street scene dataset (Cityscapes [30]), an adverse street scene dataset (ACDC [31]), an indoor understanding dataset (NYUv2 [32]), and an object-centric dataset (Pascal VOC2012 [33]). Comprehensive experiments showcase that our framework outperforms state-of-the-art KD counterparts by a large gap. Compared to the cumbersome teacher model, TransKD reduces the floating-point operations (FLOPs) by ${>}85.0\%$, while keeping competitive accuracy. By unifying patch embedding and feature map distillation, TransKD robustifies the segmentation of hard examples where previous paradigms struggle, as shown in Fig. 1. Benchmarked against the feature-map-only method Knowledge Review [10], TransKD-Base enhances the distillation performance by $5.18\%$ in mean Intersection over Union (mIoU) while adding negligible $0.21M$ parameter during the training phase, as shown in Fig. 2. On Cityscapes, TransKD improves the mIoU of the non-pretrained SegFormer-B0 [26] by $13.12\%$ and the pre-trained one by $2.09\%$. We further consider various transformer models [25, 26, 34] and verify that TransKD consistently boosts the accuracy of compact segmenters. Lastly, TransKD rivals the performance of the time-consuming pre-trained method. Our best model achieves $75.74\%$ in mIoU with only $3.72M$ parameters.
@@
  • We are the first to propose a transformer-to-transformer knowledge distillation framework in semantic segmentation and to utilize transformer-specific patch embedding as one of the knowledge sources.
  • We provide a new perspective on distilling knowledge from transformer-specific patch embeddings. Our TransKD framework unifies feature map- and patch embedding distillation.
@@
  • In-depth experiments validate the advantages of TransKD on Cityscapes, ACDC, NYUv2, and Pascal VOC2012 datasets with obvious improvements over existing knowledge distillation methods.
@@ II-A From Accurate to Efficient Semantic Segmentation
Semantic segmentation has witnessed tremendous progress since Fully Convolutional Networks (FCNs) [35] first looked at the pixel classification problem from an end-to-end perspective. Many subsequent networks followed this scheme, often raising its accuracy through encoder-decoder architectures [36] or aggregation of multi-scale context [27, 37]. An important cluster of methods leverages non-local self-attention [14, 38] to capture long-range contextual dependencies and encourage global reasoning [39, 40]. The recent success of vision transformers [13, 41] leads to the development of multiple transformer-based architectures for semantic segmentation [7, 8] which have since then become state-of-the-art on mainstream benchmarks [30, 32, 31], thanks to their ability to gather global dependencies from early layers via token mixing modules such as self-attention [14] or MLP-like blocks [42]. Yet, computational cost and lengthy training process remain a weak spot of the accuracy-driven transformer research.
On the other hand, multiple works explicitly address efficient semantic segmentation, aiming to strike a good balance between speed and accuracy. This line of research compactifies the architectures through techniques such as early downsampling [43], filter factorization [1, 15], multi-branch architecture [44, 45], and ladder-style upsampling [46, 47]. Lightweight classification backbones such as MobileNets [48] and ShuffleNets [49] have also been considered to speed up image segmentation. Moreover, SegFormer [26] introduces a lightweight and efficient semantic segmentation framework, which unifies the backbone output feature maps with a simple MLP decoder. Unlike all these works, we look at resource-efficient segmentation transformers from a new perspective of comprehensive knowledge distillation and introduce a framework for simultaneously extracting knowledge from different blocks of large-scale vision transformers and re-using it for compact and fast student transformers.
@@ II-B Vision Transformer for Dense Prediction
Transformer architecture has become the de-facto standard in Natural Language Processing (NLP) mostly due to the effectiveness of stacked self-attention and its ability to capture long-range relationships within the data. The success of self-attention-based models has also spread within the field of visual recognition, starting with the introduction of the Vision Transformer (ViT) [13] – the first transformer delivering strong results on ImageNet [11]. Subsequent developments contribute to optimizing the performance of the vision transformer, such as multi-scale networks [24, 26, 50], greater depth [51, 52], knowledge distillation [53, 54] and blending of global and local attention [22, 55].
At the same time, multiple lightweight vision transformers [54, 56] have emerged. For example, LeViT [54] is a hybrid model re-using ResNet stages within the transformer architecture while MobileViT [57] leverages MoibleNetV2 blocks [48] along the downsampling path and Mobile-Former [58] establishes a dual-stream architecture by bridging MobileNet and its transformer branch. Unfortunately, recent studies [57] suggest that MobileViT and other ViT-based models are still not efficient enough for real-time processing on mobile devices. In this work, we develop a knowledge distillation framework based on vision transformers, which consistently improves the performance of compact dense prediction and semantic segmentation transformers [26, 25, 34], leading to a much better trade-off of accuracy, computational cost, and pre-training requirements.
Figure 4: Our knowledge distillation framework TransKD. It is divided into two parts: knowledge distillation of patch embeddings (red arrows and rectangles) and feature maps (green arrows and rectangles). The loss function consists of two distillation terms (HCL and MSE) and a cross-entropy term. (a) TransKD-Base is the basic version of TransKD, constructed with two fundamental modules, CSF and PEA. (b) TransKD-GL and (c) TransKD-EA are two optimized versions of TransKD.
@@ II-C Knowledge Distillation for Semantic Segmentation
Hinton et al. [17] introduced the concept of Knowledge Distillation (KD), which learns to transfer knowledge from a large-scale model to a more compact one. These two models behave like a teacher and a student and this paradigm is widely used in computer vision [16, 59]. In this work, our goal is to study and design a KD framework specifically for semantic segmentation with transformer-based models.
KD approaches have been grouped [16] into three categories based on the form of knowledge: response-based KD [12, 9, 18], feature-based KD [9, 18, 19, 60, 61, 62], and relation-based KD [10, 20, 21]. Liu et al. [12] first introduced the idea of structured knowledge distillation, where adversarial learning is used to align the segmentation map produced by the compact student network with the one of the cumbersome teacher network. Knowledge Review (KR) [10] for the first time proposed to distill knowledge using cross-stage connection paths. Double Similarity Distillation (DSD) [20] transfers both detailed spatial dependencies and global category correlations. Unlike previous works mostly devoted to the distillation of spatial-wise knowledge, some recent works focus on the distillation of channel-wise distribution. Channel Distillation (CD) [9] focuses on the soft distributions of channels and pays attention to the most salient parts of the channel-wise maps.
Additionally, knowledge adaptation is addressed in [63] by optimizing affinity distillation. Adaptive Perspective Distillation (APD) [64] mines detailed contextual cues from each training sample, whereas [65] follows a hierarchical distillation approach and enables a one-to-all spatial matching. All of these works are dedicated to CNN-to-CNN or CNN-to-transformer knowledge distillation. In contrast, our proposed TransKD model explores this concept in a novel transformer-to-transformer fashion, by explicitly considering multi-source knowledge within the transformer architecture and taking transformer-specific patch embeddings into account.

### v4

@@ Abstract
Semantic segmentation benchmarks in the realm of autonomous driving are dominated by large pre-trained transformers, yet their widespread adoption is impeded by substantial computational costs and prolonged training durations. To lift this constraint, we look at efficient semantic segmentation from a perspective of comprehensive knowledge distillation and aim to bridge the gap between multi-source knowledge extractions and transformer-specific patch embeddings. We put forward the Transformer-based Knowledge Distillation (TransKD) framework which learns compact student transformers by distilling both feature maps and patch embeddings of large teacher transformers, bypassing the long pre-training process and reducing the FLOPs by ${>}85.0\%$. Specifically, we propose two fundamental modules to realize feature map distillation and patch embedding distillation, respectively: (1) Cross Selective Fusion (CSF) enables knowledge transfer between cross-stage features via channel attention and feature map distillation within hierarchical transformers; (2) Patch Embedding Alignment (PEA) performs dimensional transformation within the patchifying process to facilitate the patch embedding distillation. Furthermore, we introduce two optimization modules to enhance the patch embedding distillation from different perspectives: (1) Global-Local Context Mixer (GL-Mixer) extracts both global and local information of a representative embedding; (2) Embedding Assistant (EA) acts as an embedding method to seamlessly bridge teacher and student models with the teacher’s number of channels. Experiments on Cityscapes, ACDC, NYUv2, and Pascal VOC2012 datasets show that TransKD outperforms state-of-the-art distillation frameworks and rivals the time-consuming pre-training method. The source code is publicly available at https://github.com/RuipingL/TransKD.
@@ I Introduction
Semantic segmentation assigns category labels at a pixel-level (see examples in Fig. 1) and is a crucial tool in a wide range of applications, such as autonomous driving [5, 6, 7, 8] and navigation assistance for vulnerable road users [9, 10]. Driven by the achievements of deep learning, semantic segmentation has empowered precise scene comprehension in the context of autonomous driving, as evidenced by remarkable accuracies [11, 12]. Autonomous driving requires high-speed operation and the capacity to respond from significant distances to detect potential hazards. However, the computational resources available on vehicle mobile platforms are often insufficient to support large, effective semantic segmentation models, and compact but less effective models generally fail to distinguish distant details. Moreover, there has been a notable oversight in addressing the computational limitations crucial for real-world applications, including the speed of inference and training, as well as memory footprint.
While Convolutional Neural Networks (CNNs) have been the visual recognition front-runners for almost a decade, they encounter difficulties capturing large-scale interactions due to their focus on local neighborhood operations and limited receptive fields. Vision transformers [11, 12, 13] overcome this challenge through the concept of self-attention [14], which learns to intensify or weaken parts of input at a global scale. This excellent quality of modeling long-range dependencies has placed transformers at the forefront of virtually all segmentation benchmarks [11, 12, 13]. Unfortunately, transformers are usually very large and do not generalize well when trained on insufficient amounts of data due to the lack of inductive bias compared to CNNs [13], such as translation equivariance and locality, which is relevant to the semantic segmentation task. Thus, vision transformers often show less satisfactory performance on smaller task-specific datasets, and time-consuming pre-training process on a massive dataset is entailed [11, 13], e.g., ${\geq}100$ epochs on ImageNet [3]. The high computational and training costs constitute a significant bottleneck for real applications, where cumbersome models cannot be deployed on mobile devices, whereas smaller models suffer from inaccurate predictions [4, 15].
@@
A popular way to reduce the computational cost is knowledge distillation [17], which offers a mechanism for effective knowledge transfer between a large teacher model and a smaller student network and has been widely studied for deep CNNs [4, 1, 2]. Gou et al. [16] distinguish three types of knowledge to be distilled (see Fig. 3a-c): response-based knowledge [4, 1, 17], feature-based knowledge [4, 18, 19], and relation-based knowledge [2, 20, 21]. The response-based and feature-based knowledge distillation utilizes the output of a single specific layer of the teacher model (the last or the intermediate layer respectively), whereas the relation-based knowledge distillation tracks the relationship between multiple layers simultaneously. These paradigms are primarily designed with the CNN-to-CNN transfer in mind [4, 1, 2] and do not adequately cover knowledge distillation from transformer-specific intermediate blocks. The patch embedding method is a transformer-specific module to partition the input images or feature maps into a sequence of patches. It is typically achieved through convolution operations with learnable parameters, rather than by reshaping. According to the vanilla vision transformer [13], simply partitioning the input images or feature maps into large non-overlapping patches is impractical to learn the spatial relations. This limitation is addressed with optimized embedding methods, e.g. learnable positional embeddings [13], conditional positional embeddings [22, 23], shifted windows [24], and overlapping patch embeddings [25, 26]. Viewing that there is a large gap between the performances of the transformers with a simple patchifying process and optimized embedding methods, we recognize the importance of the sequential information within patch embeddings, which are processed by the embedding module that transforms feature maps into these embeddings. Therefore, we regard patch embeddings as a transformer-specific knowledge source.
@@
To achieve efficient semantic segmentation, we propose Transformer-based Knowledge Distillation (TransKD), the first framework designed to distill knowledge from large-scale semantic segmentation transformers. We focus on comprehensive distillation from different knowledge sources within the transformer-specific building blocks (our key idea is illustrated in Fig. 3d). TransKD distills the information from both (1) feature maps and (2) the transformer-specific patch embeddings. As mentioned before, patch embeddings differ from feature maps not only in the distinct shape and dimensionality but also in the inherent knowledge acquired through the patchifying process. Therefore, we consider that transformer-specific patch embeddings should be fully exploited to maximize the potential of transformer-transformer knowledge distillation. By combining knowledge from feature maps and patch embeddings, both spatial and sequential relationships can be harvested, thereby boosting efficient semantic segmentation through distillation.
@@
Specifically, we design two fundamental and two optimization modules within the TransKD framework. (1) For feature map distillation, we build a relation-based scheme atop Knowledge Review [2] and introduce the Cross Selective Fusion (CSF) module to merge cross-stage feature maps via channel attention and construct the feature map distillation streams within hierarchical transformers. (2) For patch embedding distillation, we introduce the fundamental Patch Embedding Alignment (PEA) module to perform a dimensional transformation of patch embeddings along the channel dimension to facilitate multi-stage patch embedding distillation streams. (3) Recognizing that both neighboring and long-range relationships are vital for semantic segmentation [26, 27], our Global-Local Context Mixer (GL-Mixer) extracts both global and local information of a representative embedding for distillation. (4) Furthermore, considering the large gap between the student’s and the teacher’s size that negatively impacts the knowledge distillation results [28, 29], we present the Embedding Assistant (EA) module, which acts an embedder with teacher’s number of channels. EA helps to build a pseudo teacher assistant model by combining the student’s transformer blocks and seamlessly bridge teacher and student models.
@@
We demonstrate the benefits of our approach for real-world mobile applications on four benchmarks: a general street scene dataset (Cityscapes [30]), an adverse street scene dataset (ACDC [31]), an indoor understanding dataset (NYUv2 [32]), and an object-centric dataset (Pascal VOC2012 [33]). Comprehensive experiments showcase that our framework outperforms state-of-the-art KD counterparts by a large gap. Compared to the cumbersome teacher model, TransKD reduces the floating-point operations (FLOPs) by ${>}85.0\%$, while maintaining competitive accuracy. By unifying patch embedding and feature map distillation, TransKD robustifies the segmentation of hard examples where previous paradigms struggle, as shown in Fig. 1. Benchmarked against the feature-map-only method Knowledge Review [2], TransKD-Base enhances the distillation performance by $5.18\%$ in mean Intersection over Union (mIoU) while adding negligible $0.21M$ parameter during the training phase, as shown in Fig. 2. On Cityscapes, TransKD improves the mIoU of the non-pretrained SegFormer-B0 [26] by $13.12\%$ and the pre-trained one by $2.09\%$. We further validate various transformer models [25, 26, 34] and conform that TransKD consistently boosts the accuracy of compact segmenters. Lastly, TransKD rivals the performance of the time-consuming pre-trained method. Our best model achieves $75.74\%$ in mIoU with only $3.72M$ parameters.
@@
  • We are the first to propose a transformer-to-transformer knowledge distillation framework in semantic segmentation and the first to utilize transformer-specific patch embedding as one of the knowledge sources.
  • We provide a new perspective on distilling knowledge from transformer-specific patch embeddings. Our TransKD framework unifies feature map and patch embedding distillation.
@@
  • In-depth experiments validate the advantages of TransKD on Cityscapes, ACDC, NYUv2, and Pascal VOC2012 datasets, demonstrating significant improvements over existing knowledge distillation methods.
@@ II-A From Accurate to Efficient Semantic Segmentation
Semantic segmentation has witnessed tremendous progress since Fully Convolutional Networks (FCNs) [35] first looked at the pixel classification problem from an end-to-end perspective. Many subsequent networks followed this scheme, often raising its accuracy through encoder-decoder architectures [36] or the aggregation of multi-scale context [27, 37]. An important group of methods leverages non-local self-attention [14, 38] to capture long-range contextual dependencies and promote global reasoning [39, 40]. The recent success of vision transformers [13, 41] leads to the development of multiple transformer-based architectures for semantic segmentation [11, 12] which have since then become state-of-the-art on mainstream benchmarks [30, 32, 31], thanks to their ability to capture global dependencies from early layers via token mixing modules such as self-attention [14] or Multi-Layer Perceptron (MLP) blocks [42]. Yet, the computational cost and lengthy training process remain a weak spot of the accuracy-driven transformer research.
On the other hand, multiple works explicitly address efficient semantic segmentation, aiming to strike a good balance between speed and accuracy. This line of research compactifies architectures through techniques such as early downsampling [43], filter factorization [5, 15], multi-branch architecture [44, 45], and ladder-style upsampling [46, 47]. Lightweight classification backbones such as MobileNets [48] and ShuffleNets [49] have also been adapted to expedite image segmentation. Moreover, SegFormer [26] introduces a lightweight and efficient semantic segmentation framework that unifies the backbone output feature maps with a simple MLP decoder. Unlike all these works, our TransKD views resource-efficient segmentation transformers from a new perspective of comprehensive knowledge distillation. We introduce a framework that simultaneously extracts knowledge from different blocks of large-scale vision transformers and reuses it for compact and fast student transformers.
@@ II-B Vision Transformer for Dense Prediction
Transformer architecture has become the de-facto standard in Natural Language Processing (NLP), mostly due to the effectiveness of stacked self-attention and its ability to capture long-range relationships within the data. The success of self-attention-based models has also extended to the field of visual recognition, beginning with the introduction of the Vision Transformer (ViT) [13] – the first transformer to deliver strong results on ImageNet [3]. Subsequent developments have further optimized the performance of the vision transformer, incorporating features such as multi-scale networks [24, 26, 50], increased depth [51, 52], knowledge distillation [53, 54] and a blend of global and local attention [22, 55].
At the same time, multiple lightweight vision transformers [54, 56] have emerged. For example, LeViT [54] is a hybrid model reusing ResNet stages within the transformer architecture while MobileViT [57] leverages MoibleNetV2 blocks [48] along the downsampling path and Mobile-Former [58] establishes a dual-stream architecture by bridging MobileNet and its transformer branch. Unfortunately, recent studies [57] suggest that MobileViT and other ViT-based models are still not efficient enough for real-time processing on mobile devices. In this work, we develop a knowledge distillation framework based on vision transformers, which consistently improves the performance of compact dense prediction and semantic segmentation transformers [26, 25, 34], leading to a much better trade-off of accuracy, computational cost, and pre-training requirements.
Figure 4: Our knowledge distillation framework TransKD. It is divided into two parts: knowledge distillation of patch embeddings (indicated by red arrows and rectangles) and feature maps (indicated by green arrows and rectangles). The loss function consists of two distillation terms (HCL and MSE) and a cross-entropy term. (a) TransKD-Base is the basic version of TransKD, constructed with two fundamental modules, CSF and PEA. (b) TransKD-GL and (c) TransKD-EA are two optimized versions of TransKD.
@@ II-C Knowledge Distillation for Semantic Segmentation
Hinton et al. [17] introduced the concept of Knowledge Distillation (KD), which involves transferring knowledge from a large-scale model to a more compact one. These two models act as teacher and student, and this paradigm is widely used in computer vision [16, 59]. In this work, our goal is to study and design a KD framework specifically for semantic segmentation with transformer-based models.
KD approaches have been categorized [16] into three types based on the form of knowledge: response-based KD [4, 1, 18], feature-based KD [1, 18, 19, 60, 61, 62], and relation-based KD [2, 20, 21]. Liu et al. [4] first introduced the idea of structured knowledge distillation, leveraging adversarial learning to align the segmentation map produced by the compact student network with the one of the cumbersome teacher network. Knowledge Review (KR) [2] for the first time proposed to distill knowledge using cross-stage connection paths. Double Similarity Distillation (DSD) [20] transfers both detailed spatial dependencies and global category correlations. Unlike previous works primarily devoted to the distillation of spatial-wise knowledge, some recent works focus on the distillation of channel-wise distribution. Channel Distillation (CD) [1] emphasizes the soft distributions of channels and pays attention to the most salient parts of the channel-wise maps.
Additionally, knowledge adaptation is addressed in [63] by optimizing affinity distillation. Adaptive Perspective Distillation (APD) [64] mines detailed contextual cues from each training sample, whereas [65] employs a hierarchical distillation approach and enables one-to-all spatial matching. All these works are dedicated to CNN-to-CNN or CNN-to-transformer knowledge distillation. In contrast, our proposed TransKD model explores this concept in a novel transformer-to-transformer fashion, by explicitly considering multi-source knowledge within the transformer architecture and taking transformer-specific patch embeddings into account.